OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 –3x + 2y = 1

OpenStudy (ajprincess):

what do u get when u add both the equations @homeisawesome

OpenStudy (anonymous):

how would u add them both? o.O

OpenStudy (ajprincess):

3x + y = 11-(1) –3x + 2y = 1-(2) (1)+(2)=> 3x+y-3x+2y=11+1 3y=12

OpenStudy (anonymous):

ajprincess can you help me with my question? PLease!!!

OpenStudy (ajprincess):

find the value of y. Does that help?

OpenStudy (anonymous):

y = 4? ok out of that i only get the last line..how did u get the -1 and -2?

OpenStudy (ajprincess):

ya it is right:) i jst named the equations as (1) and (2)

OpenStudy (anonymous):

oh i get it now thanks!

OpenStudy (ajprincess):

Welcome:)